# Channel Choice and Customer Value

### v1

@@ 6.2.2 Effect on Profitability
To summarize, in our empirical setting, we find that adopters due to the macro environmental shock of COVID-19 are more profitable post-adoption, relative to organic adopters despite similar spending. However, those that adopted due to firm driven promotions like Black Friday and the Loyalty Program are less profitable. An important managerial implication is that the firm should not naively assume that users who adopt online shopping due to external factors will show the same levels of profitability post-adoption relative to organic adopters. We discuss the mechanisms and theoretical underpinnings behind our findings in the next section, and we also expand on the managerial implications.
7 Mechanisms, Theoretical Underpinnings and Managerial Implications
We now discuss the mechanisms and theoretical underpinnings behind our findings on spend and profit; if and how managers can translate these findings to other settings, and some general takeaways from our analysis.
7.1 Share of offline/online spend and the theory of Consumer Inertia
One of our interesting findings from $\S$6.2 is that covid adopters are more profitable compared to organic adopters even though their overall spending is similar. We now examine if this is due to differences in the post-adoption share of offline/online spending between different groups of adopters in more detail. We also discuss the theoretical underpinnings behind these differences and derive implications for managers based on theory and our empirical findings.
Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009)would predict that adoption driven by external shocks (e.g., COVID-19, firm driven promotions) may lead to temporary disruptions in consumer habits (adopting a new channel and multichannel shopping), but without reinforcement, consumers may revert to prior routines (not use the newly adopted channels much). We test this prediction in our empirical setting, by comparing post-adoption channel usage between consumers who adopted online shopping organically and those who did so in response to the COVID-19 shock, a short-term one-time firm-driven promotion (Black Friday) and a more long-term and recurring firm driven promotion (Loyalty Program). In other words, we examine if there are differences in how the total spend with the retailer is distributed across the offline and online channels for the different groups of event adopters relative to organic adopters in the post-adoption period.
As a first step, we run the DID models specified in Equation (3) and (4) with the share of offline spend, and offline spend as the outcome variables. The results are presented in Web Appendix $\S$§E.1. We find that covid adopters spend a larger fraction of their spend offline relative to organic adopters (by about 7 percentage points) and also have a higher offline spend in the post-adoption period  [^17] Since covid adopters are more active in the offline channel, we similarly see a corresponding positive effect on the number of orders and variety of products purchased offline; see Table A26 in Web Appendix $\S$§E). Black Friday adopters also spend a larger fraction of their spending offline relative to organic adopters (by about 9 percentage points). Loyalty Program adopters also have a higher share of offline spending relative to organic adopters.
(a) Covid
(b) Black Friday
(c) Loyalty Program
From Figure 2(a), in the case of COVID-19, we see that covid adopters not only have a higher offline share of spending, post-adoption, relative to organic adopters, but their trend of moving towards more online shopping is much slower than that of organic adopters181818see Web Appendix $\S$§E.1 for formal tests of the magnitude of the difference. This is fully in-line with what consumer inertia and habit theory would predict. Further, we know from $\S$5.2 that covid adopters are older and thus are more likely to be more set in their habits and exhibit higher inertia (Yang and Ching 2014). To test whether the difference in the rate of growth of online shopping is primarily driven by differences in age, we perform a median split of all the covid adopters and organic adopters and confirm that the differences in trends only show up among older adopters and there is no difference for younger adopters [^19] See Figure A3 and Table A24 in Web Appendix E.1 for more details.
In the case of Black Friday promotion, we see from Figure 2(b) that the level of offline share of spending post-adoption is higher relative to organic adopters, again consistent with what one would expect from consumer inertia and habit theory. However, Black Friday adopters show a faster rate 202020with p-value 0.109; see Web Appendix E.1 Table A23. Extending the post-adoption period further reduces p-value. of increasing their online share of spend relative to organic adopters. This may be explained by the fact that Black Friday adopters are younger than organic adopters, and thus are less set in their habits and more open and willing to explore the conveniences of online shopping. However, since they adopted online shopping due to an external incentive, it still takes some time to catch up organic adopters in terms of post-adoption share of online spending.
Finally, in the case of the Loyalty Program, from Figure 2(c) we see again, that Loyalty Program adopters have a slightly higher share of offline spend, and there is no difference in trends relative to organic adopters. The level difference again is what we might expect from consumer inertia theory, and the fact that there is no difference in trends may be due to the fact that the two adopter groups are quite similar in terms of demographics, and also the fact that the Loyalty Program involves monthly online activation, and therefore there is some reinforcement for the habit of online shopping for Loyalty Program adopters.
Thus, based on consumer inertia theory and our empirical findings, managers should not expect consumers who adopt online shopping for external reasons to have as high a share of online spending after adoption as organic adopters. Moreover, whether the rate of growth in online share of spend for consumers who adopt online shopping due to external reasons is higher or lower relative to organic adopters depends on whether the external event/incentive is more likely to attract younger or older consumers (relative to organic adopters) to adopt online shopping [^21] For interested readers, we also explore the online-offline differences in other metrics of interest (e.g. Quantity purchased, Number of orders, Unique items, Brands, Categories, and Subcategories) in Web Appendix E.2..
7.2 Lower post-adoption spend from firm-driven promotions and the theory of forward buying
Our other interesting finding from $\S$6.2 is both groups of firm driven promotion adopters i.e. Black Friday adopters and Loyalty Program adopters exhibit lower post-adoption spend and profit relative to organic adopters. We know that promotions are an important mechanism for firms to induce various kinds of consumer behavior including the adoption of online shopping. A plausible theoretical explanation for this finding is the theory of forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017). This theory indicates that that time-limited promotions may induce forward buying, leading to demand distortion and reduced post-promotion consumption .
In our context, Black Friday adopters and Loyalty Program adopters could be increasing their spending during the promotion (i.e., the adoption period) and effectively pulling forward their future demand, resulting in a “post-promotion dip” in spending in the post-adoption period. To examine if this is the case, we first present a visual depiction of the trendlines for monthly spending of Black Friday adopters and Loyalty Program adopters relative to organic adopters and offline-only customers in Figure 3. We see that there is a spike in spend during the month of adoption, for all three types of adopters  [^22] We see a similar pattern for covid adopters as well. and the increase in spending during the month of adoption is higher for Black Friday adopters and Loyalty Program adopters relative to organic adopters.
To formally examine the differences between the two groups of event adopters relative to organic adopters during each month around adoption, we estimate the following event study regression:
We estimate the event study model in Equation (5) for Black Friday and the Loyalty Program and present the estimation results in Web Appendix §E.3. Figure 4(a) shows the point estimates and 95% confidence intervals for the dynamic treatment effects. We see a clear pattern consistent with the notion of forward buying among both Black Friday adopters and Loyalty Program adopters, relative to organic adopters. During the month of adoption, they spend more than organic adopters, and in the months following adoption, they spend less [^23] Notably, if we only look at offline spending (see Figure 4(a) in Web Appendix E.3), there are no significant differences in offline spending between Black Friday adopters and organic adopters, suggesting that the forward buying effect is specific to the online channel for Black Friday. But for the Loyalty Program, we see the effect across both channels.. Finally, we confirm 242424see Figure A5 in Web Appendix E.3. that for COVID-19, which is not a firm-driven promotion, we do not see evidence of forward buying.
Figure 3: Monthly Total Spend Visualization – Black Friday/Loyalty Program vs. Organic Adopters.
(a) Black Friday
(b) Loyalty Program
(a) Black Friday
(b) Loyalty Program
7.3 Managerial Implications
Our main findings indicate that consumers who adopt online shopping due to different reasons (organic, macro environmental shocks, or firm-driven promotional activities) behave differently in the post-adoption period. We have discussed above that our findings are explained and supported by the theories of consumer inertia and forward buying, and we believe that our main takeaways can be abstracted to other similar situations where these same theoretical concepts may apply. An important and practically useful managerial implication is that firms should not naively use the post-adoption behavior of organic adopters as a proxy to forecast the behavior of consumers who adopt online shopping for other reasons such as environmental or macroeconomic shocks, or firm-run marketing campaigns, and we highlight this implication for managers evaluating promotional campaigns to increase adoption of online shopping with some numbers from our empirical context.
To assess whether to run a promotional campaign (like the Black Friday online discounts) to induce customers to adopt online shopping, firms have to forecast the expected benefits in terms of incremental sales from consumers who choose to adopt due to the promotion, and compare the benefits to the costs of running the campaign. If managers naively use organic adopters’ incremental profit as a proxy for adopters resulting from the promotional campaign (Black Friday adopters, in this case), they could grossly overestimate the expected benefits. This in turn would lead them to wrongly compute metrics like the time taken to break-even from such a promotional campaign. To illustrate this point, we use the Black Friday promotional campaign as an example and show some back-of-the-envelope calculations.
First, we quantify the cost of running the Black Friday promotion, by computing the difference between actual consumer spending under promotional prices on Black Friday and what consumers would have paid for the same quantities at regular prices. We find the average discount across categories is around 28%, with dry food categories discounted around 20% and hygiene categories around 33%. The average promotional cost of Black Friday sales is 299.18 MCU per customer for that day. Then, from our DID results, we obtain the average monthly incremental profit of organic adopters (23.57 MCU) and Black Friday adopters (7.81 MCU)  [^25] Numbers are taken from Column 3 of Table A34 in Web Appendix $\S$§E.4.. By using the behavior of organic adopters to predict the behavior of Black Friday adopters, managers would predict that the incremental profits from Black Friday adopters would result in a break-even duration of 13 months post-adoption (i.e. 299.18 / 23.57). However, our analysis shows that it should actually take at least 39 months (e.g., 299.18 / 7.81) to break even, nearly three times longer than the managers’ initial forecast. Therefore, managers should appropriately calibrate their expectations of increased consumer spending when running promotions that encourage the adoption of online shopping.
@@ 8.1 Effect size measurement
where $w_{covid}$ is the fraction of “true” covid adopters in the sample. [^26] Assuming the monthly incidence of organic adopters remains constant, we estimate that around 38.5% of the covid adopters could actually be organic adopters (12799/33246)(12799/33246)( 12799 / 33246 ). Hence, $w_{covid}=61.5\%$ can be the fraction of “true” covid adopters. Therefore, the “true” ATT of spend for covid-driven adopters relative to organic adopters would be $\frac{-3.095}{0.615}=-5.033$ (std. error: 10.512), and the “true” ATT of profit would be $\frac{-9.079}{0.615}=14.763$ (std. error: 3.698) [^27] Numbers are taken from the column 2 of Tables 7 and 10.. This confirms that even with potential contamination, the difference in spend relative to organic adopters is small and with a confidence interval that includes zero, while the difference in profit is larger and with a confidence interval that does not include zero. As such, our substantive conclusions remain unchanged.
@@
    First, since the Black Friday sale occurs only on a single day (and not over a prolonged period, unlike COVID-19), and there is a large spike in adopters on that single day compared to the days around it, the extent of potential contamination is very small. In this case, the results are very similar to those from $\S$6.2 since historical data suggest that close to 85.7% of adopters during the Black Friday event adopt due to the event (and not organically). Nevertheless, we still conduct a similar analysis by backing out the “true” ATT using the estimated ATT of Black Friday adopters relative to organic adopters divided by $w_{black\_friday}=85.7\%$. The “true” ATT of spend for Black Friday adopters relative to organic adopters would be $\frac{-69.88}{0.857}=-81.54$ (std. error: 16.639), and the ATT of profit would be $\frac{-16.04}{0.857}=-18.716$ (std. error: 5.944). [^28] Numbers are taken from the column 4 of Tables 7 and 10. This confirms that even with such potential contamination, Black Friday adopters have lower spend and profit relative to organic adopters.
@@ 8.2 Parallel pre-trends
We test whether the parallel trends assumption holds for our DID analyses comparing spend and profitability [^29] Note that for our analyses of share of offline spend, parallel trends are satisfied by definition since all customers have 100% offline spend pre-adoption. We also present the results for our other minor metrics. between event adopters and organic adopters, and between adopters and offline-only customers, using both visual inspection and more formal event study regressions in Web Appendix $\S$§F.1. The parallel trends assumption appears to be satisfied for both spend and profit when comparing event adopters and organic adopters for all three events. However, the parallel pre-trends assumption is not satisfied for the comparisons between adopters and offline-only customers, across all three events. Further, specifically for the case of covid adopters and organic adopters, there seems to be a difference in pre-trends visually, although the differences in pre-adoption outcomes are not statistically different for most of the pre-adoption months. For these cases (spend and profits for adopters vs. offline-only customers for all three events, and covid adopters vs. organic adopters), we bound our treatment effects under different assumptions of the extent of the violation using the Honest Difference-in-Differences approach proposed by Rambachan and Roth (2023) in Web Appendix $\S$§F.2. In all instances but one, we conclude that even if the post-adoption trend violation is twice as large as the pre-adoption trend, our substantive conclusions remain unchanged in terms of both direction and statistical significance. The only exception is that for the Loyalty Program event, the difference in post-adoption profit for adopters relative to offline-only customers has a confidence interval that includes zero. However, even for this event, the difference in post-adoption spend is large and with a confidence interval that does not include zero.
@@ 8.3 Doubly robust DID, Propensity Score Matching and IPW approaches
In $\S$5.2, we saw that there are some differences in the demographics and other pre-adoption variables between organic adopters and event adopters. These differences should not affect the validity of our results as long as the parallel trends assumption holds (Angrist and Pischke 2008; Roth et al. 2023). Nevertheless, to provide additional credibility to our findings and ensure that they are robust to any factors correlated with observable characteristics that may affect the propensity to adopt online shopping organically versus due to external events, we also consider doubly robust DID (Sant’Anna and Zhao 2020) and propensity score-based approaches (Rosenbaum and Rubin 1985; Stuart et al. 2011). These methods provide better comparability between groups by improving covariate balance by discarding poorly matched customers. We compute propensity scores using a logistic regression model, with variables including demographics (e.g., gender, age, and household income) and pre-adoption purchase variables (i.e., average monthly spend, number of unique orders, unique brands, and unique categories). With the estimated propensity scores, we use our DID estimator using both IPW and PSM methods, and also run a doubly robust DID estimator. The propensity score model specification and detailed results from the IPW, PSM and doubly-robust DID methods are shown in Web Appendix $\S$§F.3. We find that our substantive conclusions remain the same with these approaches.
@@ 8.4 Alternate definitions for the organic adopter and covid adopter cohorts
To check whether our results are sensitive to the particular months we chose to define our covid adopter and organic adopter cohorts in our COVID-19 analysis, and to test whether our conclusions would change if we used “later” covid adopters vs. “early” covid adopters, we repeat our analysis with alternative definitions for both cohorts and present results in Web Appendix $\S$§F.4. To start with, we redefine organic adopters as those customers who were offline-only before 2019-10-31 and made their first online purchase between 2019-11-01 and 2020-02-29 (i.e., we extend the period we previously used to define organic adopters). We re-run the same set of DID models and find that all results are consistent with our main analysis. Next, we redefine the cohort of covid adopters as those customers who were offline-only before 2019-12-31 and made their first online purchase between 2020-05-01 and 2020-06-30 (i.e., two months later than we had used in our main analysis). Again, we find that all the results are consistent with the main analysis. Finally, we compare “early” covid adopters (who adopted online shopping between 2020-03-16 and 2020-04-30) and “later” covid adopters (who adopted online shopping between 2020-05-01 and 2020-06-30). Again, we run the same set of DID models and do not find any significant differences except that the “later” covid adopters spend slightly higher by 2.48%.303030p-value = 0.048.
@@
8.5 Using Log Scale for Dependent Variables
To test whether the positive skewness in purchase outcomes affects our results, we re-estimate all DID models using a natural log transformation for the dependent variables ($\ln(y+1)$) and report the results in Web Appendix $\S$§F.5. Across all specifications, the results remain consistent with our main findings, confirming the robustness of our analysis to log transformation of the outcomes.
8.6 Handling Zero-Inflated Dependent Variables
A potential concern in our setting is that customers do not make purchases every month, and many customer-month spending observations are zero. Therefore, we consider an alternative specification that accommodates zero-inflated dependent variable. Specifically, we estimate a two-part (hurdle) model at the customer-month level (Cameron and Trivedi 2005). This approach involves: (1) a linear probability model with a DID specification to estimate changes in the likelihood of making a purchase during the post-adoption periods [^31] Here, the dependent variable of linear probability model is $\mathbbm{I}\{Spend_{it}\}$, a binary indicator for whether customer $i$ made purchase during month $t$, and the DID specification is Equations (1) and (2) for comparing adopters vs. offline-only customers, and Equations (3) and (4) for comparing organic adopters vs. event adopters); (2) a DID model that estimates the changes in spend conditional on making a purchase. [^32] We estimate the same equations as the main analysis, but use the customer-month observations with non-zero spend. Overall, we find that the results from these specifications are consistent with our main results; see Web Appendix $\S$§F.6 for detailed estimation results.
@@ 9 Conclusions
In conclusion, this paper examines the impact of the adoption of online shopping, i.e., shopping on a retailer’s e-commerce website and/or mobile app, on customers who previously only shopped at physical stores, using data from a large pet supplies retailer in Brazil spanning the period of 2019-2024. We study four groups of consumers who adopted online shopping for different reasons during our data period: (a) covid adopters who started shopping online during the onset of COVID-19, (b) Black Friday adopters who adopted online shopping due to deep discounts available online during Black Friday, (c) Loyalty Program adopters who adopted online shopping during the launch of a loyalty program which had to be activated on the retailer’s mobile app, and (d) organic adopters who started shopping online fully of their own volition, without any external stimuli. We find that for all four types of adopters of online shopping, there is an increase in spend with the retailer post-adoption, relative to consumers who stayed offline-only.
@@
However, there are important differences between the four groups of adopters in terms of their post-adoption behavior along three dimensions - spend, profitability and share of offline/online spend. Compared to organic adopters, covid adopters, who adopted online shopping due to a macro-environmental shock, spend similarly in the post-adoption period. However, Black Friday adopters and Loyalty Program adopters, who adopted online shopping due to firm-driven promotional activities, spend less. We show that this is explained by the theory of forward buying during promotions - Black Friday adopters and Loyalty Program adopters increase their spending during the promotion, (when they adopt online shopping) by pulling forward their future demand, resulting in a “post-promotion dip” in their demand. In terms of channel usage post-adoption, consistent with consumer inertia and habit theory, we find that all three groups of externally driven adopters have a lower share of online spend post-adoption relative to organic adopters. Further, we find that covid adopters also have a slower growth in the share of online spending relative to organic adopters. This is mainly driven by the fact that covid adopters are older on average, compared to organic adopters, and thus are more set in their habits. In contrast, the younger Black Friday adopters show a faster rate of growth for online spend share relative to organic adopters. Loyalty Program adopters do not differ significantly from organic adopters in terms of the growth in the share of online spending, probably because their demographics are quite similar and the fact that consumers have to activate the loyalty program online every month. Finally, we find that covid adopters are more profitable than organic adopters due to their higher share of offline spending, where margins are higher in our setting. Black Friday adopters and Loyalty Program adopters are less profitable relative to organic adopters, due to their lower spending.
Taken together, our findings indicate that consumers who adopt online shopping due to different reasons (organic, macro environmental shocks or different types of firm driven promotional activities) behave differently post-adoption, and the differences can be explained using consumer inertia and habit theory, the theory of forward buying during promotions, and the differences in demographics between these groups of consumers. Managers can use their knowledge of consumer demographics, and the theories of consumer inertia and promotional forward buying to predict spend, profit and channel usage outcomes in other situations where these same theories apply.
An important managerial implication is that firms should not naively use the behavior of organic adopters of a new channel as a proxy to forecast the behavior of consumers who adopt the new channel for other reasons such as environmental or macroeconomic shocks, or firm-run marketing campaigns. As an example, we find that the retailer only breaks-even on running the Black Friday promotion to induce more customers to adopt online shopping, after 39 months, instead of the expected 13 months it would have predicted based on the post-adoption behavior of organic adopters. Further, even if post-adoption spending is similar (like in the case of covid adopters), differences in the usage of online channels vs. offline could cause significant differences in profitability.
Finally, we note that our study has limitations that serve as avenues for future research. First, as highlighted above, the demographics of the different types of adopters are an important input along with the theories of consumer inertia and forward buying, to predict post-adoption behavior. The external stimuli for online adoption in our setting are limited to the COVID-19 pandemic, and two specific promotional campaigns run by our retailer. Future research could investigate whether other types of events (e.g. store closures or other types of promotions or macro-shocks) lead to adopters with different demographics (and therefore potentially different interactions with these behavioral forces) than what we see in our setting. Second, we focus on one retail category (pet supplies), and it is unclear if consumers in other categories and contexts, with other forces at play, show similar behavior. For instance, recent research in the non-retail and more competitive hotel industry has found that adopters of a hotel chain’s mobile app decrease their spending at the focal hotel because they tend to also adopt competitors’ apps at the same time (Gu and Kannan 2021). Thus, future research that studies these questions in a wider set of categories and contexts can further enhance our understanding of the consequences of consumer migration to online channels.

### v2

@@ 7.2.3 Profitability
To summarize, in our empirical setting, we find that adopters due to the macro environmental shock of COVID-19 are more profitable post-adoption, relative to organic adopters despite similar spending. However, those who adopted due to firm-driven promotions like Black Friday and the Loyalty Program are less profitable. An important managerial implication is that the firm should not naively assume that users who adopt online shopping due to external factors will show the same levels of profitability post-adoption relative to organic adopters.
8 Managerial Implications
Our main findings indicate that consumers who adopt online shopping through different channel pathways (organic, macro-environmental shocks, or promotions) behave differently in the post-adoption period. These patterns are well explained by consumer inertia and forward buying, and should generalize to other settings where similar mechanisms operate. In particular, forward buying and stockpiling can reduce the long-run spending increase from promotion-induced online adopters. To mitigate this and maximize long-run value, firms should design incentives that discourage stockpiling (e.g., phased or behavior-contingent rewards) and combine promotions with rapid follow-up nudges to encourage repeat (rather than one-off) behavior. For consumers pushed online by shocks (like COVID-19), habit-reinforcing promotions and personalized nudges in the immediate post-adoption window can help convert trials of online channels into stable habits, thereby improving long-run profitability. Beyond these general lessons, we derive concrete implications for two common marketing tasks: (a) CLV forecasting and (b) ROI/breakeven analysis for promotions.
8.1 Implications for CLV forecasts
Customer Lifetime Value (CLV) is a core metric for customer relationship management, targeting, and budgeting (Fader, Hardie, and Lee 2005; Gupta et al. 2006; Fader, Hardie, and Shang 2010). Our results show that when offline-only customers adopt online shopping for different reasons (e.g., Black Friday, COVID-19, loyalty program, or organically), their spending and profitability trajectories differ markedly by pathway. Ignoring this heterogeneity leads to biased forecasts of incremental CLV from online adoption and, in turn, to suboptimal decisions, for example, overestimating the effectiveness of promotions aimed at inducing online adoption.
We quantify the average incremental (relative to offline-only customers) CLV from online adoption for each adopter cohort $s$. Following standard practice (Gupta et al. 2006), the (average) incremental CLV equals the present value of all future incremental monthly profits attributable to adoption under an infinite horizon:
$Incremental~CLV_{s}=\sum_{t=1}^{\infty}\frac{ATT_{s}\times r_{s}^{t}}{(1+i)^{t}}=ATT_{s}\frac{r_{s}}{1+i-r_{s}},$ (6)
where: (1) $ATT_{s}$ is the monthly average incremental profit (our ATT estimate) for cohort $s$ relative to offline-only customers; (2) $i$ is the firm’s monthly cost of capital (WACC); and (3) $r_{s}$ is the (average) monthly retention rate of cohort $s$ post-adoption, defined as the probability of purchasing next month conditional on purchasing in the focal month. [^22] We follow Gupta, Lehmann, and Stuart (2004) and use an infinite-horizon formulation to avoid arbitrary truncation; our results are qualitatively similar under a finite horizon. All quantities are monthly. The $ATT_{s}$ estimates come from our DID models for each event, and retention rates are obtained non-parametrically. We report monthly retention rates for organic adopters and event adopters for each event in Table 8. Based on discussions with the retailer, we set WACC to 15% annually, implying a monthly cost of capital of $i=(1+15\%)^{1/12}-1=1.17\%$.
Table 8 presents the incremental CLV from online adoption for organic versus event adopters within each event (COVID-19, Black Friday, Loyalty Program). Covid adopters have an incremental CLV of 73.486 MCU, 36.4% higher than organic adopters (i.e., (73.486 - 53.859) / 53.859). By contrast, promotion-induced adopters are far less valuable in the long run: Black Friday adopters’ incremental CLV is 69.2% lower, and Loyalty Program adopters’ is 78.9% lower than that of organic adopters. These differences mirror our monthly profit findings: covid adopters generate higher long-run value, while promotion-induced adopters (Black Friday, Loyalty Program) generate substantially lower value than organic adopters.
Table 8: Incremental CLV (measured in MCU) of Organic Adopters and Event Adopters – By Three Events (COVID-19, Black Friday, and Loyalty Program)
COVID-19 Black Friday Loyalty Program
Organic Adopters
ATT 30.280 23.570 18.550
Monthly Retention Rate 0.648 0.638 0.559
Incremental CLV 53.859 40.319 22.873
Event Adopters
ATT 39.780 7.813 2.986
Monthly Retention Rate 0.656 0.621 0.625
Incremental CLV 73.486 12.417 4.831
Diff in Incremental CLV of Event Adopters vs. organic adopters +36.4% -69.2% -78.9%
[…]
We next show how CLV forecasts go wrong when managers ignore adoption pathways. We consider two common scenarios:
  • $\bullet$
    Using organic adopters as proxies. Suppose the firm launches a new promotion to encourage online adoption and proxies long-term CLV using organic adopters (because those are the only adopters observed pre-promotion). Table 8 shows that doing so overstates the incremental CLV of promotion-induced adopters by 69.2% for Black Friday and 78.9% for the loyalty program. Using organic adopters as proxies can therefore severely overvalue promotion-induced adopters.
  • $\bullet$
    Pooled forecasts. Alternatively, the firm might pool event and organic adopters and estimate a single incremental CLV for all adopters. The results of this exercise are shown in Table 9.232323For the COVID-19 event, for example, the pooled ATT from Table 4 is 37.30 MCU, and the pooled monthly retention rate is 0.654. Using Equation (6), the incremental CLV is $37.30\times\frac{0.654}{1+0.0117-0.654}=68.332$ MCU. We perform similar calculations for Black Friday and the loyalty program events. Pooled estimates overstate the value of promotion-induced adopters by almost 60% (Black Friday) and 80% (Loyalty Program), while understating the value of COVID-induced adopters by 7.5%.
Table 9: Incremental CLV (measured in MCU) Under Pooled Estimation – By Three Events (COVID-19, Black Friday, and Loyalty Program)
COVID-19 Black Friday Loyalty Program
ATT (Pooled) 37.370 18.440 18.410
Monthly Retention Rate (Pooled) 0.654 0.633 0.565
Incremental CLV (Pooled) 68.332 30.822 23.301
Diff in Incremental CLV of Event Adopters vs. Pooled Estimates +7.5% -59.7% -79.3%
[…]
In both scenarios, failing to distinguish adoption pathways biases CLV forecasts—typically overstating the value of promotion-induced adopters and making weak promotional strategies appear attractive, which can lead to overspending and erode long-run profitability. A key takeaway is that retailers should (a) estimate (incremental) CLV separately by adoption pathway and (b) set pathway-specific caps on the cost of inducing online adoption, aligning offer depth and paid-media spend with those caps.
8.2 Implications for ROI and Breakeven Analyses for Promotions
To assess whether to run a promotional campaign (such as Black Friday online discounts or an app-activated loyalty program) to induce online adoption, firms must compare the expected incremental profits from adopters to the cost of the campaign. If managers naïvely use organic adopters’ incremental profit as a proxy for promotion-induced adopters, they will overstate expected benefits and miscalculate breakeven times and ROI. We illustrate this with the Black Friday and Loyalty Program campaigns.
[…]
In sum, managers must calibrate expectations of incremental spending when running promotions to encourage online adoption. Ignoring adoption pathways can lead to overly optimistic ROI assessments and underestimation of the time required to recoup promotional investments.
@@ 9.1 Effect size measurement
[…]
@@
[…]
@@ 9.2 Parallel pre-trends
[…]
@@ 9.3 Doubly robust DID, Propensity Score Matching and IPW approaches
[…]
@@ 9.4 Alternate definitions for the organic adopter and covid adopter cohorts
[…]
@@
9.5 Alternate operationalizations of dependent variables
[…]
@@ 10 Conclusions
This paper examines how the adoption of a retailer’s online channels by previously offline-only customers affects their subsequent behavior, and how this effect varies by the reason for adoption. Building on prior work on multichannel shopping, and the theories of forward buying, and consumer inertia, we argue that (i) adopting online channels and becoming multichannel should increase overall spending relative to remaining offline-only, but (ii) the magnitude, persistence, and channel mix of this increase should differ systematically across adoption pathways. Using transaction data from a large Brazilian pet supplies retailer (2019–2024), we study four adopter groups: covid adopters (adoption during the onset of COVID-19), Black Friday adopters (promotion-driven adoption via deep online discounts), Loyalty Program adopters (adoption triggered by a loyalty program activated on the app), and organic adopters (voluntary adoption without identifiable external stimuli).
@@
Consistent with our predictions, all four adopter types increase their spending relative to offline-only customers. However, their post-adoption trajectories differ markedly. Relative to organic adopters, covid adopters exhibit similar total spend but, in line with forward-buying theory, Black Friday adopters and Loyalty Program adopters spend less in the post-adoption period, having stocked up during the promotion. In terms of channel mix, and consistent with consumer inertia and habit theory, all three externally driven adopter groups have a lower online spend share than organic adopters. Moreover, covid adopters show slower growth in online share, reflecting their older and more habit-driven profile, while younger Black Friday adopters catch up more quickly. Loyalty Program adopters do not differ significantly from organic adopters in the growth of online share, likely because of similar demographics and the monthly online activation requirement, which reinforces online use. Finally, consistent with typical cross-channel margin differences, covid adopters are more profitable than organic adopters due to their higher offline spend share, whereas Black Friday adopters and Loyalty Program adopters are less profitable due to lower overall spending.
These results highlight that the motives and circumstances surrounding online adoption play a crucial role in shaping long-term multichannel behavior, through mechanisms linked to demographics, forward buying, and consumer inertia. Managerially, our findings underscore the need to tailor incentives to the adoption context. Promotions designed to induce online adoption should be structured to limit stockpiling (e.g., via phased or behavior-contingent rewards) and followed by targeted nudges to sustain engagement. For customers who move online due to external shocks, habit-reinforcing incentives can help convert temporary adoption into lasting behavior, improving long-run profitability. We also show that forecasts of incremental Customer Lifetime Value (CLV) and breakeven analyses can be substantially biased if adoption pathways are ignored: using organic adopters as proxies or pooling all adopters can overstate the value of promotion-induced adopters and mislead ROI assessments. Firms should therefore explicitly distinguish organic from externally driven adopters when forecasting CLV and evaluating promotions.
Our study has limitations that suggest directions for future research. First, the external stimuli in our setting are limited to one macro shock (COVID-19) and two specific promotions run by a single retailer. Future work could examine other events (e.g., store closures, different promotion designs, or alternative macro shocks) that may attract different demographic profiles and interact differently with forward buying and inertia. Second, we focus on one retail category (pet supplies) in a single country. Studying similar questions in other categories and contexts would help assess the generalizability of our findings. Finally, we estimate descriptive ATTs of online adoption using a difference-in-differences approach with consumer and time fixed effects, without fully separating self-selection from causal effects of adoption. We do so because the overall ATT, which embeds both selection and causal components, is the quantity most relevant for managerial decisions. Nonetheless, future academic work could complement our analysis by estimating a causal ATE of online adoption, for example using instrumental variables or other identification strategies.